OpenStudy (randomecorner):

What value of k makes the equation true? 12 + (–17) = –17 + k k = pleas help :3

OpenStudy (anonymous):

k=12

OpenStudy (randomecorner):

how does it though>

OpenStudy (randomecorner):

oh ok wow i feel dumb thx

OpenStudy (anonymous):

Let's solve your equation step-by-step. 1217+k Step 1: Simplify both sides of the equation. 5=k17 step 2: Flip the equation. k17=5 Step 3: Add 17 to both sides. k17+17=5+17 k=12

OpenStudy (anonymous):

Your not dumb you just need a little extra help every one needs a little help

OpenStudy (anonymous):

And np

OpenStudy (randomecorner):

thank you :p

OpenStudy (randomecorner):

is it ok if i right this down for notes?